Langley Corporation currently collects all of its customer payments in Greensboro. By going to a new lockbox system with boxes in Chicago and Boulder, Langley expects to reduce the time from when customers mail their checks to when the funds are collected. It expects the time to fall from an average of 9 days to an average of 4 days, a 5-day savings. Langley collects $50,000 per day. The extra costs associated with the lockbox are $15,000 per year. Langley's opportunity cost of funds is 10% per year. What is the expected annual before-tax profit of using the new lockbox system?
Answer #1

Reduction in cash balance due to lock box = number of days saved*average cash collection per day = 5*50,000 = 250,000

Return earned on this cash = opportunity cost of funds*cash = 10%*250,000 = 25,000

Annual before-tax profit = return earned - annual cost of lock box = 25,000 - 15,000 = 10,000